physiological body temperature in humans is around 37 degree . At this temperature, the ion product constant for water, Kw, is 2.4x10^-14. What is the ph of pure, neutral water at 37 degree
Answer #1

In pure water the hydrogen ion concentration must be equal to the hydroxide ion concentration .

Kw= 2.4×10‐¹⁴ = [H+]×[OH-]

= [H+]×[H+]

= [H+]²

So [H+]= 1.55×10-⁷

So pH= -log[H+]

= 6.81 (ANSWER)